OpenStudy (anonymous):

Help! Factor this expression as a difference of squares. (Medal for step by step.) a^2+4a+4-b^2

OpenStudy (anonymous):

I need help too! can you get me more followers :/ I only have like 200

OpenStudy (anonymous):

I'll meet you half way :D

OpenStudy (anonymous):

Help 4 Help.

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

@xiphoidx could you help meout with my gradpoint im on 2b and it is way to hard

OpenStudy (anonymous):

a 2 +4a+4=(a+2) 2 giving (a+2) 2 −b 2 and you get get (a+2+b)(a+2−b) because the difference of squares formula is a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

it depends on what it is

OpenStudy (anonymous):

..What? Where'd the (a+2) 2 and the other numbers come from?

OpenStudy (anonymous):

i just started 2b and it is so hard like could you plz help me

OpenStudy (anonymous):

a+2)^2 is also a^2+4a+4 if you foil (a+2)(a+2) and then add in that b^2 I forgot to make them exponents